There are 2 women participants in a chess tournament . the number of men played between themselves exceeded by 52 the number of game they played with the women .if each player played one game with each other, find  the number of men in the tournament 

Grade:11

1 Answers

AKASH GOYAL AskiitiansExpert-IITD
13 years ago

Dear Rahul

Let us assume no. of men =n

no. of games played between themselve =nC2

no. games men played with women=2n

According to question

nC2-2n=52

n*(n-1)/2-2n=52

n*n-5n-104=0

After solving

n=13 or -8

-8 is not possible

So no. of men =13

All the best

AKASH GOYAL
AskiitiansExpert-IITD

Please feel free to post as many doubts on our discussion forum as you can. We are all IITians and here to help you in your IIT JEE preparation.


Now you can win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.